OpenStudy (anonymous):

A right triangle has legs measuring 8 and 3.5. What is the measure of the smallest angle?

OpenStudy (anonymous):

That depends. In either case, "3.5" is the shortest side, and will be opposite to the smallest angle (SA). If 8 is the hypotenuse, then sin(SA) = 3.5/8 Which means SA = sin^-1(3.5/8) If 8 is NOT the hypotenuse, then it must be adjacent to the smallest angle which means tan(SA) = 3.5/8 Which means SA = tan^-1(3.5/8)